# Tunable absorption spectrum splitting in a pulse-driven three-level system

### v1

@@ I Introduction
Beyond the two-level systems, when a third energy level is included, for a three-level system e.g. in the $\Lambda$ configuration, a driving field can give rise to electromagnetically induced transparency [12, 13] with prominent applications in quantum memories [14, 15, 16, 17, 18]. The parent phenomenon, the Autler-Townes splitting, is observed for a strongly driven system and has also been implemented as a protocol for quantum memories [19, 20, 21, 22, 23]. Given the aforementioned effects of pulse sequences on two-level systems, it is then natural to investigate their effects on three-level systems and to explore potential advantages that they can provide in this context. In the present paper, we study the absorption spectrum of the three-level system in the ladder or cascade configuration when the transition between the bottom two levels is driven by a periodic pulse sequence while the transition between the middle and the highest energy level is probed with a weak field. We find that the resulting absorption spectrum displays similarities with the Autler-Townes doublet with separation between the main peaks defined by the inverse of the inter-pulse delay. In addition, this spectrum shows little dependence on the pulse carrier frequency. These results demonstrate a flexible capacity to modulate the absorption spectrum of a three-level system with achievable pulse protocols, highlighting alternative implementations of quantum memories in these systems.
@@ II Model and Numerical Solution
We consider a three-level system in the ladder configuration with energy levels $|1\rangle$, $|2\rangle$, and $|3\rangle$ with respective energies, $E_{1}$, $E_{2}$ and $E_{3}$ with $E_{1}<E_{2}<E_{3}$. This system is placed in a bosonic bath of photons to which it is coupled. In the usual formulation of the Autler-Townes effect, (Fig. 1 (a)) the $|1\rangle\leftrightarrow|2\rangle$ transition is driven strongly while the $|2\rangle\leftrightarrow|3\rangle$ transition is gauged by a weak probing field. The measured absorption spectrum displays, when the driving field is resonant with the $|1\rangle\leftrightarrow|2\rangle$ transition, two symmetric peaks around the transition frequency that are separated by the Rabi frequency ($\Omega_{1}$) of the driving field. This splitting is due to the dressed states, around the ground and the excited states, that are formed in the resonantly driven two-level system. A finite detuning ($\Delta_{1}$) between the driving field frequency $\omega_{1}$ and the transition frequency leads to an asymmetry in the position and the magnitude of the peaks.
Here, we consider the situation where the $|1\rangle\leftrightarrow|2\rangle$ transition is instead driven by a periodic sequence of finite width $\pi_{x}$ pulses (Fig. 1 (b)). To this end, we apply a driving field with a time dependent Rabi frequency $\Omega_{1}(t)$ such that, $\Omega_{1}(t)=\Omega_{1}$ during the pulse and is zero otherwise. Similarly to the conventional Autler-Townes splitting setup, the probing field, at frequency $\omega=E_{3}-E_{2}+\Delta_{2}$ , with a Rabi frequency $\Omega_{2}\ll\Omega_{1}$, is applied at all times. The driving field is finite for time $t_{\pi}=\pi/\Omega_{1}$. The relaxation rate from state $|2\rangle$ to state $|1\rangle$ is $\Gamma_{2}$ while the relaxation rate from state $|3\rangle$ to state $|2\rangle$ is $\Gamma_{3}$, and there is no direct transition allowed between states $|1\rangle$ and $|3\rangle$. We set $\Gamma_{2}=2$ and we measure our energy and time in units of $\Gamma_{2}$ and $1/\Gamma_{2}$ respectively.
@@
Where $H_{A}$, $H_{F}$, and $H_{AF}$ are, respectively, the Hamiltonian describing the atom, the field modes and the coupling between the atom and the field mode. $H_{d}(t)$ corresponds to the driving field on the $|1\rangle\leftrightarrow|2\rangle$ transition. While $H_{p}(t)$ describes the probing field coupled to the $|2\rangle\leftrightarrow|3\rangle$ transition.
The driving field is treated at a semi-classical level which is consistent with driving the emitter with a coherent field [3].
To characterize the time evolution of the system, we will use the density matrix operator of the 3-level system when the bosonic degrees of freedom are integrated out of the problem. This density matrix operator of the 3-level system is given by
$\rho$ $=$ $\rho_{11}|1\rangle\langle 1|+\rho_{22}|2\rangle\langle 2|+\rho_{33}|3\rangle\langle 3|$ (2)
$+$ $\rho_{21}|2\rangle\langle 1|+\rho_{12}|1\rangle\langle 2|+\rho_{32}|3\rangle\langle 2|$
$+$ $\rho_{23}|2\rangle\langle 3|+\rho_{31}|3\rangle\langle 1|+\rho_{13}|1\rangle\langle 3|$
@@
$\dot{\rho}_{11}$ $=$ $\Gamma_{2}\rho_{22}-i\frac{\Omega_{1}}{2}\rho_{21}+i\frac{\Omega_{1}}{2}\rho^{*}_{21}$ (3)
$\dot{\rho}_{22}$ $=$ $-\Gamma_{2}\rho_{22}+\Gamma_{3}\rho_{33}+i\frac{\Omega_{1}}{2}\rho_{21}$ (4)
-- $i\frac{\Omega_{1}}{2}\rho^{*}_{21}-i\frac{\Omega_{2}}{2}\rho_{32}+i\frac{\Omega_{2}}{2}\rho^{*}_{32}$
@@
$\dot{\rho}_{21}$ $=$ $-i\frac{\Omega_{1}}{2}\rho_{11}+i\frac{\Omega_{1}}{2}\rho_{22}$ (6)
@@
$+$ $\left[i\Delta_{2}-\frac{1}{2}\Gamma_{32}\right]\rho_{32}+i\frac{\Omega_{1}}{2}\rho_{31}$
$\dot{\rho}_{31}$ $=$ $-i\frac{\Omega_{2}}{2}\rho_{21}+i\frac{\Omega_{1}}{2}\rho_{32}$ (8)
@@
Here factors of $\mathrm{e}^{i\omega_{1}t}$, $\mathrm{e}^{i\omega t}$, and $\mathrm{e}^{i(\omega_{1}+\omega)t}$ have been included in $\rho_{21}$, $\rho_{32}$ and $\rho_{31}$ respectively. Note also that $\rho_{12}$ and $\rho_{23}$ have respectively been replaced in these equations by $\rho_{21}^{*}$ and $\rho_{32}^{*}$, using the above-mentioned properties of the density matrix operator. $\Gamma_{2}$ and $\Gamma_{3}$ are the transition rates between the $|2\rangle$ and the $|1\rangle$ states, and between the $|3\rangle$ and the $|2\rangle$ states respectively. In addition, we have introduced the parameter $\Gamma_{32}=\Gamma_{3}+\Gamma_{2}$. For a continuously driven $|1\rangle\leftrightarrow|2\rangle$ transition, a steady expression for the occupation of the highest energy level has been derived [24]. For the pulse driven system, however, the solution becomes untractable and we resort to a numerical solution.
@@
At time $t=0$, the system is prepared in the state $|2\rangle$. All matrix elements are set to zero except for $\rho_{22}=1$. The transition between the lower energy level $|1\rangle$ and the intermediate energy level $|2\rangle$ is then driven periodically with inter-pulse delay $\tau$, and pulses applied by a field of Rabi frequency $\Omega_{1}$, that is applied for a time $\pi/\Omega_{1}$ amounting to a finite width $\pi$-pulse. The weak probing field, with Rabi frequency $\Omega_{2}(\ll\Omega_{1})$ is applied at all times on the transition between the intermediate energy level $|2\rangle$ and the upper energy level $|3\rangle$.
@@
We integrate these differential equations iteratively starting from the initial conditions for the density matrix operator $\rho(t_{0}=0)$ and calculate $\rho(t_{0}+\delta t)$, $\rho(t_{0}+2\delta t)$ up to $\rho(t_{0}+N_{t}\delta t)=\rho(t_{max})$. Where $t_{max}=N_{t}\delta t$ and $N_{t}$ is the number of time steps of width $\delta t$. The time step $\delta t$ is chosen such that the discretization error is negligible. The time evolution of the density matrix characterized by equations (3 - 8), is carried out for frequencies $\omega$ of the weak probing field, covering a range of values of $\Delta_{2}$ from $\omega_{min}=-40$ to $\omega_{max}=40$ outside of which there is negligible spectral weight. For each probing field frequency, we evolve the system up to a maximum time when the system has reached the steady state. The steady state here is identified by tracking the time evolution of the density matrix operator, particularly the occupation number of the highest excited state, $\rho_{33}$, and confirming that it is no longer changing appreciably.
The absorption spectrum of interest is then defined by the occupation of the highest excited state, $\rho_{33}$, in this steady state, for varying values $\omega$ of the probing field frequency.
@@
Figure 2: Population of the second excited state (|3)|3$\rangle$)| 3 ⟩ ) as a function of time, $\rho_{33}(t)$, when the $(|1\rangle\leftrightarrow|2\rangle)$ transition is driven by a periodic sequence of $\pi$ pulses with period $\tau=0.3$. The pulse carrier frequency is resonant with the transition, $\Delta_{1}=0$, and the pulses are due to a driving field of Rabi frequency $\Omega_{1}=50$ that is periodically applied for time $t_{\pi}=\pi/\Omega$. $\rho_{33}(t)$ is shown for $\omega=0$ (solid green line), $\pi/2\tau$ (dashed blue line) and $3\pi/2\tau$ (dashed red line).
@@
Figure 3: Population of the second excited state or absorption spectrum of the $(|2\rangle\leftrightarrow|3\rangle)$ transition when the $(|1\rangle\leftrightarrow|2\rangle)$ is driven by a periodic sequence of $\pi$ pulses with period $\tau$ the pulse is due to a driving field of Rabi frequency $\Omega_{1}=50$ that is applied for time $t_{\pi}=\pi/\Omega_{1}$. Absorption spectrum shown for detuning $\Delta_{1}=0$ under pulse sequences of different periods: $\tau=0.1$ (black solid line), $\tau=0.2$ (dashed blue line), $\tau=0.3$ (dashed green line), $\tau=0.4$ (dotted red line).

### v2

@@ I Introduction
Beyond the two-level systems, when a third energy level is included, for a three-level system e.g. in the $\Lambda$ configuration, a driving field can give rise to electromagnetically induced transparency [12, 13] with prominent applications in quantum memories [14, 15, 16, 17, 18]. The parent phenomenon, the Autler-Townes splitting, is observed for a strongly driven system and has also been implemented as a protocol for quantum memories [19, 20, 21, 22, 23]. Given the aforementioned effects of pulse sequences on two-level systems, it is then natural to investigate their effects on three-level systems and to explore potential advantages that they can provide in this context. In the present paper, we study the absorption spectrum of the three-level system in the ladder or cascade configuration when the transition between the bottom two levels is driven by a periodic pulse sequence with inter-pulse delay $\tau$ while the transition between the middle and the highest energy level is probed with a weak field. We find that the resulting absorption spectrum displays similarities with the Autler-Townes doublet with separation between the main peaks defined by the inverse of the inter-pulse delay. In addition, this spectrum shows little dependence on the pulse carrier frequency. These results demonstrate a flexible capacity to modulate the absorption spectrum of a three-level system with achievable pulse protocols, highlighting alternative implementations of quantum memories in these systems. The pulse sequence produces an effect analogous to that of the continuously driven lower two levels with a Rabi frequency equal to $\pi/\tau$ while probing the transition between the first excited state and the highest excited state by a weak probing field. The findings in this paper are consistent with previous theoretical and experimental studies of pulse-induced tuning of spectral signatures of two and three-level systems[24, 25, 26].
@@ II Model and Numerical Solution
We consider a three-level system in the ladder configuration with energy levels $|1\rangle$, $|2\rangle$, and $|3\rangle$ with respective energies, $E_{1}$, $E_{2}$ and $E_{3}$ with $E_{1}<E_{2}<E_{3}$. This system is placed in a bosonic bath of photons to which it is coupled. In the usual formulation of the Autler-Townes effect, (Fig. 1 (a)) the $|1\rangle\leftrightarrow|2\rangle$ transition is driven strongly while the $|2\rangle\leftrightarrow|3\rangle$ transition is gauged by a weak probing field. The measured absorption spectrum displays, when the driving field is resonant with the $|1\rangle\leftrightarrow|2\rangle$ transition, two symmetric peaks around the transition frequency that are separated by the Rabi frequency ($\Omega_{1}$) of the driving field. This splitting is due to the dressed states, around the ground and the excited states, that are formed in the resonantly driven two-level system. A finite detuning ($\Delta_{1}$) between the driving field frequency $\nu_{1}$ and the transition frequency leads to an asymmetry in the position and the magnitude of the peaks.
Here, we consider the situation where the $|1\rangle\leftrightarrow|2\rangle$ transition is instead driven by a periodic sequence of finite width $\pi_{x}$ pulses (Fig. 1 (b)). To this end, we apply a driving field with a time dependent Rabi frequency $\Omega_{d}(t)$ such that, $\Omega_{d}(t)=\Omega_{1}$ during the pulse and is zero otherwise. Similarly to the conventional Autler-Townes splitting setup, the probing field, at frequency $\hbar\omega=E_{3}-E_{2}+\Delta_{2}$ , with a Rabi frequency $\Omega_{2}\ll\Omega_{1}$, is applied at all times. The driving field is non-zero for time $t_{\pi}=\pi/\Omega_{1}$. The relaxation rate from state $|2\rangle$ to state $|1\rangle$ is $\Gamma_{2}$ while the relaxation rate from state $|3\rangle$ to state $|2\rangle$ is $\Gamma_{3}$, and there is no direct transition allowed between states $|1\rangle$ and $|3\rangle$. We set $\Gamma_{2}=2$ and we measure our energy and time in units of $\Gamma_{2}/2$ and $2/\Gamma_{2}$ respectively.
@@
Where the individual terms of the Hamiltonian describe, respectively, the atomic energy levels ($H_{A}$), the radiation field modes ($H_{F}$), the coupling between the atom and the field mode ($H_{AF}$), the driving field on the $|1\rangle\leftrightarrow|2\rangle$ transition ($H_{d}(t)$), the probing field coupled to the $|2\rangle\leftrightarrow|3\rangle$ transition $($$H_{p}(t)$)). These individual terms of the Hamiltonian in Eq.(1) are defined by:
$H_{A}$ $=$ $\hbar\omega_{1}|1\rangle\langle 1|+\hbar\omega_{2}|2\rangle\langle 2|+\hbar\omega_{3}|3\rangle\langle 3|,$ (2)
$H_{F}$ $=$ $\sum_{k}\hbar\omega_{k}a_{k}^{\dagger}a_{k},$ (3)
$H_{AF}$ $=$ $i\sum_{k}g_{12}^{k}a^{\dagger}_{k}|1\rangle\langle 2|+i\sum_{k}g_{13}^{k}a^{\dagger}_{k}|1\rangle\langle 3|$ (4)
$\;\;\;\;\;\;\;\;\;\;\;\;+i\sum_{k}g_{23}^{k}a^{\dagger}_{k}|2\rangle\langle 3|+H.c.$
$H_{d}$ $=$ $\frac{\Omega_{d}(t)}{2}\mathrm{e}^{-i\nu_{1}t}|1\rangle\langle 2|+H.c.,$ (5)
$H_{p}$ $=$ $\frac{\Omega_{2}}{2}\mathrm{e}^{-i\omega t}|2\rangle\langle 3|+H.c.,$ (6)
$g^{k}_{12},\;g^{k}_{13},\;g^{k}_{23}$ are, respectively, the coupling strengths of the transition between state $|1\rangle$ and state $|2\rangle$, state $|1\rangle$ and state $|3\rangle$, and, state $|2\rangle$ and state $|3\rangle$ with the radiation mode $k$. it is assumed that $g^{k}_{23}$ is negligible for all modes. The operators $a^{\dagger}_{k}$ and $a_{k}$ are, respectively, the creation and the annihilation operators for the $k$ radiation mode with frequency $\omega_{k}$. The relaxation rate $\Gamma_{2}$ of the $|2\rangle\to|1\rangle$ transition is given by $\Gamma_{2}=2\pi\int(g^{k}_{12})^{2}\;\delta(\omega_{k}-(\omega_{2}-\omega_{1}))\;dk$ and similarly for $\Gamma_{3}$. The driving field and the probing field are treated at a semi-classical level which is consistent with driving the emitter with a coherent field [3].
To characterize the time evolution of the system, we will use the density matrix operator of the 3-level system when the bosonic degrees of freedom are integrated out of the problem. This density matrix operator of the 3-level system is given by: $\rho=\rho_{11}|1\rangle\langle 1|+\rho_{22}|2\rangle\langle 2|+\rho_{33}|3\rangle\langle 3|+\rho_{21}|2\rangle\langle 1|+\rho_{12}|1\rangle\langle 2|+\rho_{32}|3\rangle\langle 2|+\rho_{23}|2\rangle\langle 3|+\rho_{31}|3\rangle\langle 1|+\rho_{13}|1\rangle\langle 3|$.
@@
$\dot{\rho}_{11}$ $=$ $\Gamma_{2}\rho_{22}-i\frac{\Omega_{d}(t)}{2}\rho_{21}+i\frac{\Omega_{d}(t)}{2}\rho^{*}_{21}$ (7)
$\dot{\rho}_{22}$ $=$ $-\Gamma_{2}\rho_{22}+\Gamma_{3}\rho_{33}+i\frac{\Omega_{d}(t)}{2}\rho_{21}$ (8)
- $i\frac{\Omega_{d}(t)}{2}\rho^{*}_{21}-i\frac{\Omega_{2}}{2}\rho_{32}+i\frac{\Omega_{2}}{2}\rho^{*}_{32}$
@@
$\dot{\rho}_{21}$ $=$ $-i\frac{\Omega_{d}(t)}{2}\rho_{11}+i\frac{\Omega_{d}(t)}{2}\rho_{22}$ (10)
@@
$+$ $\left[i\Delta_{2}-\frac{1}{2}\Gamma_{32}\right]\rho_{32}+i\frac{\Omega_{d}(t)}{2}\rho_{31}$
$\dot{\rho}_{31}$ $=$ $-i\frac{\Omega_{2}}{2}\rho_{21}+i\frac{\Omega_{d}(t)}{2}\rho_{32}$ (12)
@@
Here factors of $\mathrm{e}^{i\nu_{1}t}$, $\mathrm{e}^{i\omega t}$, and $\mathrm{e}^{i(\nu_{1}+\omega)t}$ have been included in $\rho_{21}$, $\rho_{32}$ and $\rho_{31}$ respectively. Note also that $\rho_{12}$ and $\rho_{23}$ have respectively been replaced in these equations by $\rho_{21}^{*}$ and $\rho_{32}^{*}$, using the above-mentioned properties of the density matrix operator. We have introduced the parameter $\Gamma_{32}=\Gamma_{3}+\Gamma_{2}$. For a continuously driven $|1\rangle\leftrightarrow|2\rangle$ transition, a steady expression for the occupation of the highest energy level has been derived [27]. For the pulse driven system, however, the solution becomes untractable and we resort to a numerical solution.
@@
At time $t=0$, the system is prepared in the state $|2\rangle$. All matrix elements of the density matrix operator are set to zero except for $\rho_{22}=1$. The transition between the lower energy level $|1\rangle$ and the intermediate energy level $|2\rangle$ is then driven periodically with inter-pulse delay $\tau$, and pulses applied by a field of Rabi frequency $\Omega_{1}$, that is applied for a time $\pi/\Omega_{1}$ amounting to a finite width $\pi$-pulse. More explicitly, the evolution of the master equation (7-12) is carried out periodically with a $\tau$ time interval corresponding, first to an evolution with $\Omega_{d}(t)=0$ for time $\tau-t_{\pi}$, followed, for time $t_{\pi}$ to an evolution with $\Omega_{d}(t)=\Omega_{1}$. The weak probing field, with Rabi frequency $\Omega_{2}(\ll\Omega_{1})$ is applied at all times on the transition between the intermediate energy level $|2\rangle$ and the upper energy level $|3\rangle$.
@@
We integrate these differential equations iteratively starting from the initial conditions for the density matrix operator $\rho(t_{0}=0)$ and calculate $\rho(t_{0}+\delta t)$, $\rho(t_{0}+2\delta t)$, … up to $\rho(t_{0}+N_{t}\delta t)=\rho(t_{max})$. Where $t_{max}=N_{t}\delta t$ and $N_{t}$ is the number of time steps of width $\delta t$. The time step $\delta t$ is chosen such that the discretization error is negligible. The time evolution of the density matrix characterized by equations (7 - 12), is carried out for frequencies $\omega$ of the weak probing field, covering a range of values of $\Delta_{2}$ from $\omega_{min}=-40$ to $\omega_{max}=40$ outside of which there is negligible spectral weight. For each probing field frequency, we evolve the system up to a maximum time when the system has reached the steady state. The steady state here is identified by tracking the time evolution of the density matrix operator, particularly the occupation number of the highest excited state, $\rho_{33}$, and confirming that it is no longer changing appreciably.
@@
Figure 2: Population of the second excited state ($|3\rangle)$ as a function of time, $\rho_{33}(t)$, when the $(|1\rangle\leftrightarrow|2\rangle)$ transition is driven by a periodic sequence of $\pi$ pulses with period $\tau=0.3$. The pulse carrier frequency is resonant with the transition, $\Delta_{1}=0$, and the pulses are due to a driving field of Rabi frequency $\Omega_{1}=50$ that is periodically applied for time $t_{\pi}=\pi/\Omega$. $\rho_{33}(t)$ is shown for $\omega=0$ (solid green line), $\pi/2\tau$ (dashed blue line) and $3\pi/2\tau$ (dashed red line). Time is measured in units of $2/\Gamma_{2}$
The absorption of the probing field, of interest in the present study, is proportional to the population of the energy level $|3\rangle$. Thus, it is studied through the occupation of the highest excited state, $\rho_{33}$, in this steady state, for varying values $\omega$ of the probing field frequency. The population of the highest energy level is probed within our weak probe treatment and in the steady state. The lineshape in this long-time regime is found, within the weak probe regime, to not depend on the actual value of the probing field Rabi frequency as long as it does not significantly modify the dynamics. However, the observed population of the highest energy level, in its magnitude, may have some dependence on the probing field Rabi frequency.
@@
Figure 3: Population of the second excited state or absorption spectrum of the $(|2\rangle\leftrightarrow|3\rangle)$ transition when the $(|1\rangle\leftrightarrow|2\rangle)$ is driven by a periodic sequence of $\pi$ pulses with period $\tau$ the pulse is due to a driving field of Rabi frequency $\Omega_{1}=50$ that is applied for time $t_{\pi}=\pi/\Omega_{1}$. Absorption spectrum shown for detuning $\Delta_{1}=0$ under pulse sequences of different periods: $\tau=0.1$ (black dashed line), $\tau=0.2$ (dashed blue line), $\tau=0.3$ (dashed green line), $\tau=0.4$ (dotted red line). The solid lines show the continuously driven system with frequency $\pi/\tau$ and all other parameters identical to the pulse-driven case. All energies are measured in units of $\Gamma_{2}/2$.
Figure 4: Separation between the main peaks as a function of $1/\tau$ for $\Delta_{1}=0$. The red line and symbols represent the data while the dashed blue line represents the linear fit with slope $\pi$ confirming the dependence on $\tau$. All energies are measured in units of $\Gamma_{2}/2$.
Figure 5: Population of the second excited state or absorption spectrum of the $(|2\rangle\leftrightarrow|3\rangle)$ transition when the $(|1\rangle\leftrightarrow|2\rangle)$ is driven by a periodic sequence of imperfect pulses accomplishing a $0.8\pi$ rotation (dashed green line), $0.9\pi$ rotation (dashed red line), $1.0\pi$ rotation (solid black line), $1.1\pi$ rotation (dashed blue line), $1.2\pi$ rotation (dashed purple line), with period $\tau=0.3$ and $\Omega_{1}=50$. The detuning with the pulse carrier frequency is $\Delta_{1}=0$. All energies are measured in units of $\Gamma_{2}/2$.
Figure 6: Population of the second excited state or absorption spectrum of the $(|2\rangle\leftrightarrow|3\rangle)$ transition when the $(|1\rangle\leftrightarrow|2\rangle)$ is driven by a periodic sequence of $\pi$ pulses with period $\tau$ the pulse is due to a driving field of Rabi frequency $\Omega_{1}=100$ (solid back line), $\Omega_{1}=50$ (solid red line), $\Omega_{1}=25$ (dashed blue line), that is applied for time $t_{\pi}=\pi/\Omega_{1}$. The pulse period is $\tau=0.2$ and the detuning is $\Delta_{1}=0$ (top) and $\Delta_{1}=3$ (bottom). All energies are measured in units of $\Gamma_{2}/2$.
Figure 7: Population of the second excited state or absorption spectrum of the $(|2\rangle\leftrightarrow|3\rangle)$ transition when the $(|1\rangle\leftrightarrow|2\rangle)$ is driven by a periodic sequence of $\pi$ pulses with period $\tau=0.2$ (top) and $\tau=0.3$ (bottom). The $\pi$ pulses are achieved with Rabi frequency $\Omega_{1}=50$. The detuning with the pulse carrier frequency is $\Delta_{1}=0$ (solid black line), $\Delta_{1}=3$ (dashed blue line), $\Delta_{1}=5$ (dashed green line), $\Delta_{1}=8$ (dotted red line). All energies are measured in units of $\Gamma_{2}/2$.